# Global AI Governance Overview: Understanding Regulatory Requirements Across Global Jurisdictions

### v1

@@ 3.4 EU Penalties
It is important to note that the Code of Practice itself does not establish or impose any penalties or sanctions. The GPAI Code of Practice is a voluntary tool, prepared by independent experts in a multi-stakeholder process, designed to help industry comply with the AI Act’s obligations for providers of general-purpose AI models. The Commission has indicated that it can favourably take adherence to the Code into account when assessing compliance with the AI Act, which may have an effect in deciding upon the amount of regulatory fines. However, all penalty provisions and enforcement mechanisms derive exclusively from the EU AI Act itself, particularly Articles 99-101. The Code of Practice serves only as a compliance demonstration tool and does not create additional legal obligations or sanctions beyond those already established in the AI Act. The penalties are defined by each member state separately, which means that the fines may differ from country to country. Article 99 AI Act (8) [euaiact2024, Article 99]: Each Member State shall lay down rules on to what extent administrative fines may be imposed on public authorities and bodies established in that Member State.
@@
  • Non-compliance with the prohibition of the AI practices referred to in Article 5 [euaiact2024, Article 5] shall be subject to administrative fines of up to EUR 35 000 000 or, if the offender is an undertaking, up to 7% of its total worldwide annual turnover for the preceding financial year, whichever is higher [euaiact2024, Article 99(3)];
  • Non-compliance with any of the following provisions related to operators or notified bodies, other than those laid down in Article 5 [euaiact2024, Article 5] , shall be subject to administrative fines of up to EUR 15 000 000 or, if the offender is an undertaking, up to 3% of its total worldwide annual turnover for the preceding financial year, whichever is higher [euaiact2024, Article 99(4)];
  • The supply of incorrect, incomplete or misleading information to notified bodies or national competent authorities in reply to a request shall be subject to administrative fines of up to EUR 7 500 000 or, if the offender is an undertaking, up to 1% of its total worldwide annual turnover for the preceding financial year, whichever is higher [euaiact2024, Article 99(5)];
  • In the case of Small and Medium-sized Enterprises (hereinafter SMEs), including start-ups, each fine referred to in this Article shall be up to the percentages or amount referred to in paragraphs 3, 4 and 5, whichever thereof is lower [euaiact2024, Article 99(6)].
@@ 3.5 Challenges caused by intersection of GDPR, DSA, and AI Act
    For example, in 2022, YouTube’s AI mistakenly removed a series of videos documenting human rights abuses in Ukraine, classifying them as violent content [barrister2024freespeech].
@@ 3.6 The US Regulatory Base
While the EU has established comprehensive enforcement mechanisms, the regulatory landscape in the United States presents a different approach. The US hasn’t established a legal framework yet. The president Trump has announced an AI Plan which is directed at ”establishing AI dominance in the world”. As we can notice, the US’ approach differs from EU, because the US is aiming at implementing less restrictions and controls. Certain regulations concerning AI are established only in specific states, with the highest number of laws in California, but on the Federal level there are no regulations yet. As of June 2025, more than half of U.S. states have enacted at least one AI or algorithmic accountability law. In total, 48 states and Puerto Rico have introduced AI-related bills this year, with 26 enacting new measures. Common priorities include transparency, child protection, and reducing algorithmic bias [beckage2025aigovernance]. These laws, including amendments to the California Consumer Privacy Act (CCPA) and specific requirements for AI training data, generative AI (gen AI) disclosures and content labeling, impose new compliance obligations that span a range of sectors, demanding significant operational and technological adjustments.
@@ 3.6.1 California
California AI laws apply based on user location, not company registration. Following the same extraterritorial approach as GDPR and CCPA, these laws regulate any company that makes AI systems accessible to California residents, regardless of where the company is incorporated, headquartered, or operates from. For example, SB 1001 [california_sb1001_2018] applies to communications with ”persons in California,” while SB 942 [california_sb942_2024] targets AI systems that are ”publicly accessible in California” with over 1 million monthly users. This user-centric jurisdictional approach means that any AI company serving California residents, whether based in Silicon Valley, New York, London, or Tokyo, must comply with California’s AI regulations, creating a ”California Effect” similar to how GDPR applies globally to companies processing EU residents’ data [pillsbury2024california].
@@ 3.6.2 Utah
Utah’s SB 149 [utah_sb149_2024] (2024) creates the Artificial Intelligence Policy Act, establishing a comprehensive regulatory framework for AI use in the state. The law establishes liability for AI use that violates consumer protection laws if not properly disclosed, requires clear disclosure when individuals interact with AI in regulated occupations or consumer transactions, and creates the Office of Artificial Intelligence Policy within the Department of Commerce to oversee AI regulation and policy development. The legislation establishes an AI Learning Laboratory Program to assess technologies, risks, and policy implications, while enabling temporary mitigation of regulatory impacts during AI pilot testing through regulatory agreements. The law also clarifies that using AI to commit offenses doesn’t exempt someone from criminal responsibility, with enforcement handled by the Division of Consumer Protection through fines up to \$2,500 per violation and court remedies. Notably, the AI Policy Act is set to repeal on May 1, making this essentially a pilot program designed to test regulatory approaches while balancing AI innovation with consumer protection through a ”sandbox” framework that allows experimentation under state oversight.
@@ 3.6.3 Arkansas
Arkansas HB 1958 [arkansas_hb1958_2025] requires all public entities in the state to create comprehensive policies regarding the authorized use of artificial intelligence and automated decision tools, with the critical requirement that a human employee or designee always makes the final decision regardless of AI recommendations. The legislation mandates that public entities develop training programs for employees on appropriate AI use, create disciplinary procedures for policy violations, and make these policies publicly available upon request. The policies must include prohibitions on using technology resources for personal political opinions, illegal activities, or circumventing security procedures, establishing a framework for responsible and controlled AI use within Arkansas’s public sector while ensuring human oversight remains paramount in government decision-making processes.
Arkansas Act 927 [arkansas_act927_2025] establishes that individuals who provide prompts or data to generative AI tools will generally own the resulting content or trained model, provided the input data is legally obtained and the content doesn’t infringe on existing copyrights or intellectual property rights. The law includes a work-for-hire provision specifying that when employees use generative AI tools as part of their job duties under employer direction and control, the employer owns the resulting content and model training data. This legislation makes Arkansas one of the first U.S. states to define ownership of generative AI output and training results, providing crucial legal clarity for individuals and businesses using AI tools while emphasizing that ownership rights cannot override existing intellectual property protections.
@@ 3.6.4 Maryland
Maryland HB 956 [maryland_hb956_2025] establishes the Workgroup on Artificial Intelligence Implementation to monitor and make recommendations related to AI regulation, consumer protection, current private sector AI use, enforcement authority for the Attorney General’s Office of Consumer Protection, and impact on government benefits determination. The workgroup will examine AI’s impact on individuals’ life opportunities, labor and employment concerns, privacy rights, consumer disclosure standards, and enforcement mechanisms, with required annual reports to legislative committees starting July 1, 2026 [maryland_hb956_2025]. Signed by Governor Wes Moore on April 22, 2025 [maryland_hb956_2025], the legislation reflects Maryland’s proactive approach to studying AI regulation before implementing comprehensive policies, though it faced criticism from consumer advocates who argued the workgroup composition favored industry representatives over civil society voices.
@@ 3.6.5 Montana
Montana SB 212 Right to Compute Act [montana_sb212_2025] establishes the nation’s first ”Right to Compute Act,” affirming citizens’ fundamental right to own, access, and use computational resources including hardware, software, and AI tools under the state constitution’s guarantees of property and free expression. The legislation requires deployers of critical infrastructure facilities controlled by AI systems to develop risk management policies based on recognized national and international standards, such as the National Institute of Standards and Technology’s AI Risk Management Framework, with annual policy reviews and fallback mechanisms to ensure human control can be restored swiftly. Signed into law by Governor Greg Gianforte on April 17, 2025, the act imposes strict limitations on government restrictions over computational rights, requiring any limitations to be demonstrably necessary and narrowly tailored to fulfill compelling public health or safety interests, while positioning Montana as a leader in protecting individual digital freedoms and technological innovation.
@@ 3.6.6 West Virginia
West Virginia HB 3187 [westvirginia_hb3187_2025] amends the existing AI Task Force established within the Office of the Governor to add identification of economic opportunities to its agenda and require annual reports to the legislature. The task force, composed of state agency heads, legislators, and representatives from AI, cybersecurity, and health industries, is responsible for recommending AI definitions for legislation, assessing AI’s impact on workforce and employment, and developing public sector best practices. Approved by the governor on April 25, 2025, and effective July 9, 2025, the legislation extends the task force’s termination date to July 1, 2027, positioning West Virginia to systematically explore AI’s potential for economic development while ensuring responsible implementation in state government operations.
@@ 3.6.7 Illinois
Illinois HB 3773 [illinois_hb3773_2025] amends the Illinois Human Rights Act to explicitly prohibit employers from using artificial intelligence that subjects employees to discrimination based on protected characteristics, including the specific prohibition against using ZIP codes as a proxy for protected classes. Effective January 1, 2026, the law requires employers to notify employees when AI is used for recruitment, hiring, promotion, training selection, discharge, discipline, tenure, or other employment decisions, though specific notice requirements will be developed by the Illinois Department of Human Rights through future rulemaking. The legislation defines AI broadly as any machine-based system that generates predictions, recommendations, or decisions, including generative AI, and establishes civil rights violations for discriminatory AI use with enforcement through the existing Human Rights Commission process, making Illinois the second state after Colorado to address algorithmic discrimination in employment decisions.
@@ 3.6.8 NIST AI RMF
The National Institute of Standards and Technology (NIST), a non-regulatory federal agency within the U.S. Department of Commerce, serves as the primary standards-setting body for the United States government and plays a crucial role in developing technological standards and guidelines. While NIST is a U.S. government agency, its frameworks and standards are designed as voluntary guidance that can be adopted by organizations worldwide, carrying no legislative authority or binding legal requirements. In response to growing concerns about artificial intelligence risks and the need for systematic risk management approaches, NIST developed the Artificial Intelligence Risk Management Framework (AI RMF 1.0). This comprehensive framework provides organizations with a structured methodology for identifying, assessing, and mitigating risks associated with AI systems throughout their lifecycle, establishing foundational principles for trustworthy AI development and deployment [nist2023airmf].
@@ 3.7.1 ISO
Other useful frameworks are brought by International Organization for Standardization (hereinafter ISO) and the key ones that can help govern AI are ISO/IEC 42005 [iso42005:2025], ISO/IEC 42001 [iso42001:2023], ISO/IEC 23894 [iso23894:2023]. These ISO standards operate as voluntary international frameworks rather than mandatory federal regulations, providing organizations with structured approaches to AI governance that can be implemented regardless of jurisdiction. These standards offer practical guidance on AI risk management, quality management systems, and ethical considerations that can serve as foundational building blocks for regulatory compliance. For example, for AI Act compliance specifically, these ISO frameworks can be particularly valuable because the EU regulation explicitly recognizes harmonized European standards as a pathway to demonstrate conformity, and ISO standards often inform the development of these harmonized standards. Organizations that proactively implement ISO AI governance frameworks can establish robust documentation, risk assessment procedures, and quality management systems that align with AI Act requirements for transparency, accountability, and risk mitigation, potentially streamlining their compliance efforts and providing evidence of due diligence to regulators across multiple jurisdictions. Let’s consider the difference and main purposes of each of these frameworks:
@@ 3.8 Asia-Pacific Regulatory Base
The Asian AI regulatory landscape presents a diverse and evolving approach to artificial intelligence governance, reflecting the region’s position as a global leader in AI development and innovation. While the Asia-Pacific region is expected to become the world’s largest AI market by 2030, with countries like China, Japan, Singapore, and South Korea driving significant technological advancement, regulatory approaches vary considerably across jurisdictions [xenoss2025apac].
@@ 3.8.1 China
All AI platforms must register AI services, undergo security reviews, label AI-generated content, and ensure data and foundation models come from legitimate, rights-respecting sources. Service providers are also held accountable for content created through their platforms. This framework builds on years of groundwork and ties into broader laws like the Personal Information Protection Law [PIPL2021] and network data security regulations [xenoss2025apac].
Administrative Provisions on Algorithmic Recommendation Services [china2022algorithmic]
@@
Provisions on the Administration of Deep Synthesis Internet Information Services [china2022deepfakes]
@@
Interim Measures for the Management of Generative AI Services [china2023generativeai]
@@ 3.8.2 Japan
Key documents include the Social Principles of Human-Centric AI (2019) [cabinet2019humancentricAI], which outline Japan’s vision for ethical, inclusive, and dignity-centered AI; the AI Governance Guidelines for Business [meti2024aiguidelinesV1], most recently updated in April 2024, offering practical tools for risk management, transparency, and human oversight; and the AI Governance Framework (2024) [grjapan2024aiJapan], which consolidates earlier recommendations into a single reference point for voluntary compliance and systematic risk assessment. These soft-law instruments are reinforced by existing statutes such as the Act on the Protection of Personal Information (APPI) — Japan’s GDPR-equivalent — the Digital Platform Transparency Act, which targets fairness in e-commerce and advertising, and the Copyright Act, relevant to generative AI outputs.

### v2

@@ 3.4 EU Penalties
It is important to note that the Code of Practice itself does not establish or impose any penalties or sanctions. The GPAI Code of Practice is a voluntary tool, prepared by independent experts in a multi-stakeholder process, designed to help industry comply with the AI Act’s obligations for providers of general-purpose AI models. The Commission has indicated that it can favourably take adherence to the Code into account when assessing compliance with the AI Act, which may have an effect in deciding upon the amount of regulatory fines. However, all penalty provisions and enforcement mechanisms derive exclusively from the EU AI Act itself, particularly Articles 99-101. The Code of Practice serves only as a compliance demonstration tool and does not create additional legal obligations or sanctions beyond those already established in the AI Act. The penalties are defined by each member state separately, which means that the fines may differ from country to country. Article 99 AI Act (8) [18, Article 99]: Each Member State shall lay down rules on to what extent administrative fines may be imposed on public authorities and bodies established in that Member State.
@@
  • Non-compliance with the prohibition of the AI practices referred to in Article 5 [18, Article 5] shall be subject to administrative fines of up to EUR 35 000 000 or, if the offender is an undertaking, up to 7% of its total worldwide annual turnover for the preceding financial year, whichever is higher [18, Article 99(3)];
  • Non-compliance with any of the following provisions related to operators or notified bodies, other than those laid down in Article 5 [18, Article 5] , shall be subject to administrative fines of up to EUR 15 000 000 or, if the offender is an undertaking, up to 3% of its total worldwide annual turnover for the preceding financial year, whichever is higher [18, Article 99(4)];
  • The supply of incorrect, incomplete or misleading information to notified bodies or national competent authorities in reply to a request shall be subject to administrative fines of up to EUR 7 500 000 or, if the offender is an undertaking, up to 1% of its total worldwide annual turnover for the preceding financial year, whichever is higher [18, Article 99(5)];
  • In the case of Small and Medium-sized Enterprises (hereinafter SMEs), including start-ups, each fine referred to in this Article shall be up to the percentages or amount referred to in paragraphs 3, 4 and 5, whichever thereof is lower [18, Article 99(6)].
@@ 3.5 Challenges caused by intersection of GDPR, DSA, and AI Act
    For example, in 2022, YouTube’s AI mistakenly removed a series of videos documenting human rights abuses in Ukraine, classifying them as violent content [39].
@@ 3.6 The US Regulatory Base
While the EU has established comprehensive enforcement mechanisms, the regulatory landscape in the United States presents a different approach. The US hasn’t established a legal framework yet. The president Trump has announced an AI Plan which is directed at ”establishing AI dominance in the world”. As we can notice, the US’ approach differs from EU, because the US is aiming at implementing less restrictions and controls. Certain regulations concerning AI are established only in specific states, with the highest number of laws in California, but on the Federal level there are no regulations yet. As of June 2025, more than half of U.S. states have enacted at least one AI or algorithmic accountability law. In total, 48 states and Puerto Rico have introduced AI-related bills this year, with 26 enacting new measures. Common priorities include transparency, child protection, and reducing algorithmic bias [40]. These laws, including amendments to the California Consumer Privacy Act (CCPA) and specific requirements for AI training data, generative AI (gen AI) disclosures and content labeling, impose new compliance obligations that span a range of sectors, demanding significant operational and technological adjustments.
@@ 3.6.1 California
California AI laws apply based on user location, not company registration. Following the same extraterritorial approach as GDPR and CCPA, these laws regulate any company that makes AI systems accessible to California residents, regardless of where the company is incorporated, headquartered, or operates from. For example, SB 1001 [4] applies to communications with ”persons in California,” while SB 942 [5] targets AI systems that are ”publicly accessible in California” with over 1 million monthly users. This user-centric jurisdictional approach means that any AI company serving California residents, whether based in Silicon Valley, New York, London, or Tokyo, must comply with California’s AI regulations, creating a ”California Effect” similar to how GDPR applies globally to companies processing EU residents’ data [38].
@@ 3.6.2 Utah
Utah’s SB 149 [42] (2024) creates the Artificial Intelligence Policy Act, establishing a comprehensive regulatory framework for AI use in the state. The law establishes liability for AI use that violates consumer protection laws if not properly disclosed, requires clear disclosure when individuals interact with AI in regulated occupations or consumer transactions, and creates the Office of Artificial Intelligence Policy within the Department of Commerce to oversee AI regulation and policy development. The legislation establishes an AI Learning Laboratory Program to assess technologies, risks, and policy implications, while enabling temporary mitigation of regulatory impacts during AI pilot testing through regulatory agreements. The law also clarifies that using AI to commit offenses doesn’t exempt someone from criminal responsibility, with enforcement handled by the Division of Consumer Protection through fines up to \$2,500 per violation and court remedies. Notably, the AI Policy Act is set to repeal on May 1, making this essentially a pilot program designed to test regulatory approaches while balancing AI innovation with consumer protection through a ”sandbox” framework that allows experimentation under state oversight.
@@ 3.6.3 Arkansas
Arkansas HB 1958 [2] requires all public entities in the state to create comprehensive policies regarding the authorized use of artificial intelligence and automated decision tools, with the critical requirement that a human employee or designee always makes the final decision regardless of AI recommendations. The legislation mandates that public entities develop training programs for employees on appropriate AI use, create disciplinary procedures for policy violations, and make these policies publicly available upon request. The policies must include prohibitions on using technology resources for personal political opinions, illegal activities, or circumventing security procedures, establishing a framework for responsible and controlled AI use within Arkansas’s public sector while ensuring human oversight remains paramount in government decision-making processes.
Arkansas Act 927 [1] establishes that individuals who provide prompts or data to generative AI tools will generally own the resulting content or trained model, provided the input data is legally obtained and the content doesn’t infringe on existing copyrights or intellectual property rights. The law includes a work-for-hire provision specifying that when employees use generative AI tools as part of their job duties under employer direction and control, the employer owns the resulting content and model training data. This legislation makes Arkansas one of the first U.S. states to define ownership of generative AI output and training results, providing crucial legal clarity for individuals and businesses using AI tools while emphasizing that ownership rights cannot override existing intellectual property protections.
@@ 3.6.4 Maryland
Maryland HB 956 [32] establishes the Workgroup on Artificial Intelligence Implementation to monitor and make recommendations related to AI regulation, consumer protection, current private sector AI use, enforcement authority for the Attorney General’s Office of Consumer Protection, and impact on government benefits determination. The workgroup will examine AI’s impact on individuals’ life opportunities, labor and employment concerns, privacy rights, consumer disclosure standards, and enforcement mechanisms, with required annual reports to legislative committees starting July 1, 2026 [32]. Signed by Governor Wes Moore on April 22, 2025 [32], the legislation reflects Maryland’s proactive approach to studying AI regulation before implementing comprehensive policies, though it faced criticism from consumer advocates who argued the workgroup composition favored industry representatives over civil society voices.
@@ 3.6.5 Montana
Montana SB 212 Right to Compute Act [35] establishes the nation’s first ”Right to Compute Act,” affirming citizens’ fundamental right to own, access, and use computational resources including hardware, software, and AI tools under the state constitution’s guarantees of property and free expression. The legislation requires deployers of critical infrastructure facilities controlled by AI systems to develop risk management policies based on recognized national and international standards, such as the National Institute of Standards and Technology’s AI Risk Management Framework, with annual policy reviews and fallback mechanisms to ensure human control can be restored swiftly. Signed into law by Governor Greg Gianforte on April 17, 2025, the act imposes strict limitations on government restrictions over computational rights, requiring any limitations to be demonstrably necessary and narrowly tailored to fulfill compelling public health or safety interests, while positioning Montana as a leader in protecting individual digital freedoms and technological innovation.
@@ 3.6.6 West Virginia
West Virginia HB 3187 [43] amends the existing AI Task Force established within the Office of the Governor to add identification of economic opportunities to its agenda and require annual reports to the legislature. The task force, composed of state agency heads, legislators, and representatives from AI, cybersecurity, and health industries, is responsible for recommending AI definitions for legislation, assessing AI’s impact on workforce and employment, and developing public sector best practices. Approved by the governor on April 25, 2025, and effective July 9, 2025, the legislation extends the task force’s termination date to July 1, 2027, positioning West Virginia to systematically explore AI’s potential for economic development while ensuring responsible implementation in state government operations.
@@ 3.6.7 Illinois
Illinois HB 3773 [24] amends the Illinois Human Rights Act to explicitly prohibit employers from using artificial intelligence that subjects employees to discrimination based on protected characteristics, including the specific prohibition against using ZIP codes as a proxy for protected classes. Effective January 1, 2026, the law requires employers to notify employees when AI is used for recruitment, hiring, promotion, training selection, discharge, discipline, tenure, or other employment decisions, though specific notice requirements will be developed by the Illinois Department of Human Rights through future rulemaking. The legislation defines AI broadly as any machine-based system that generates predictions, recommendations, or decisions, including generative AI, and establishes civil rights violations for discriminatory AI use with enforcement through the existing Human Rights Commission process, making Illinois the second state after Colorado to address algorithmic discrimination in employment decisions.
@@ 3.6.8 NIST AI RMF
The National Institute of Standards and Technology (NIST), a non-regulatory federal agency within the U.S. Department of Commerce, serves as the primary standards-setting body for the United States government and plays a crucial role in developing technological standards and guidelines. While NIST is a U.S. government agency, its frameworks and standards are designed as voluntary guidance that can be adopted by organizations worldwide, carrying no legislative authority or binding legal requirements. In response to growing concerns about artificial intelligence risks and the need for systematic risk management approaches, NIST developed the Artificial Intelligence Risk Management Framework (AI RMF 1.0). This comprehensive framework provides organizations with a structured methodology for identifying, assessing, and mitigating risks associated with AI systems throughout their lifecycle, establishing foundational principles for trustworthy AI development and deployment [36].
@@ 3.7.1 ISO
Other useful frameworks are brought by International Organization for Standardization (hereinafter ISO) and the key ones that can help govern AI are ISO/IEC 42005 [28], ISO/IEC 42001 [26], ISO/IEC 23894 [25]. These ISO standards operate as voluntary international frameworks rather than mandatory federal regulations, providing organizations with structured approaches to AI governance that can be implemented regardless of jurisdiction. These standards offer practical guidance on AI risk management, quality management systems, and ethical considerations that can serve as foundational building blocks for regulatory compliance. For example, for AI Act compliance specifically, these ISO frameworks can be particularly valuable because the EU regulation explicitly recognizes harmonized European standards as a pathway to demonstrate conformity, and ISO standards often inform the development of these harmonized standards. Organizations that proactively implement ISO AI governance frameworks can establish robust documentation, risk assessment procedures, and quality management systems that align with AI Act requirements for transparency, accountability, and risk mitigation, potentially streamlining their compliance efforts and providing evidence of due diligence to regulators across multiple jurisdictions. Let’s consider the difference and main purposes of each of these frameworks:
@@ 3.8 Asia-Pacific Regulatory Base
The Asian AI regulatory landscape presents a diverse and evolving approach to artificial intelligence governance, reflecting the region’s position as a global leader in AI development and innovation. While the Asia-Pacific region is expected to become the world’s largest AI market by 2030, with countries like China, Japan, Singapore, and South Korea driving significant technological advancement, regulatory approaches vary considerably across jurisdictions [44].
@@ 3.8.1 China
All AI platforms must register AI services, undergo security reviews, label AI-generated content, and ensure data and foundation models come from legitimate, rights-respecting sources. Service providers are also held accountable for content created through their platforms. This framework builds on years of groundwork and ties into broader laws like the Personal Information Protection Law [9] and network data security regulations [44].
Administrative Provisions on Algorithmic Recommendation Services [19]
@@
Provisions on the Administration of Deep Synthesis Internet Information Services [20]
@@
Interim Measures for the Management of Generative AI Services [21]
@@ 3.8.2 Japan
Key documents include the Social Principles of Human-Centric AI (2019) [3], which outline Japan’s vision for ethical, inclusive, and dignity-centered AI; the AI Governance Guidelines for Business [33], most recently updated in April 2024, offering practical tools for risk management, transparency, and human oversight; and the AI Governance Framework (2024) [22], which consolidates earlier recommendations into a single reference point for voluntary compliance and systematic risk assessment. These soft-law instruments are reinforced by existing statutes such as the Act on the Protection of Personal Information (APPI) — Japan’s GDPR-equivalent — the Digital Platform Transparency Act, which targets fairness in e-commerce and advertising, and the Copyright Act, relevant to generative AI outputs.